OpenStudy (anonymous):

I don't get what it mean by ''real numbers a and b'' solve the following equation for the real numbers a and b. ((1+i)/(1-i))^2+(1/(a+ib))=1+i

OpenStudy (amistre64):

thats because you are a victim of bad naming ...

OpenStudy (amistre64):

real numbers are numbers found on the real number line ...

OpenStudy (anonymous):

Complex numbers come in the form a+bi so they want you to simplify the expression to the form a+bi

OpenStudy (anonymous):

so they just want to simplify it right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

it is just saying solve for a and b but since you are working with complex numbers they are specifying that a and b are real numbers and not say \[a=x+iy\] for example

OpenStudy (anonymous):

they could have written \[(\frac{1-i}{1+i})^2+\frac{1}{z}=1+i\] solve for z